OpenStudy (anonymous):

Marble C, with mass 5.0 g, moves at a speed of 20.0 cm/s. It colides with asecond marble,D, with a mass of 10.0 g, moving at 10 cm/s in the same direction. After the collison, marble C continues with a speed of 8.0 cm/s in the same direction. -Sketch the system and identify the system. Identify the before and after situations and set up a coordinate system.

OpenStudy (anonymous):

5*20 + 10*10 = 8 *5 = 10*x, here x = Vd2 Linear momentum conservation above . To the right is positive. Mc* Vc1 + Md*Vd1 = Mc*Vc2 + Md*Vd2 Vd2 = 16m/s to the right. You can sketch this right? Just balls labeled C and D in a line, all always moving to the right with speeds indicated in the states above. You can name your y coordinate axes in any which way shape or form (orthogonal to the x) because it won't affect the x movement of the marbles if you are on a horizontal plane. Hope that helps!

OpenStudy (anonymous):

Okay, I understand how to sketch the before part, just im not certain about the after part

OpenStudy (anonymous):

ok one sec i'll draw!

OpenStudy (anonymous):

oops also i meant vd / 16 CM/s, not m/s.

OpenStudy (anonymous):

Okay, thank you